# Dual Consolidation for Pre-Trained Model-Based Domain-Incremental Learning

### v1

@@ B.1 Dataset introduction
  • DomainNet Peng et al. (2019) [^1] https://ai.bu.edu/M3SDA/ is a dataset of common objects in six different domains. All domains include 345 categories of objects, such as Bracelets, planes, birds, and cellos. The domains include clipart — a collection of clipart images; real — photos and real-world images; sketch — sketches of specific objects; infograph — infographic images with specific objects; painting — artistic depictions of objects in the form of paintings, and quickdraw — drawings of the worldwide players of the game ‘Quick Draw!’. We use the officially recommended version ‘Cleaned’ in this paper.
  • Office-Home Venkateswara et al. (2017) [^2] https://hemanthdv.github.io/officehome-dataset/ is a benchmark dataset for domain adaptation which contains four domains where each domain consists of 65 categories. The four domains are Art — artistic images in the form of sketches, paintings, ornamentation, etc.; Clipart — a collection of clipart images; Product — images of objects without a background; and Real-World — images of objects captured with a regular camera. It contains 15,500 images, with an average of around 70 images per class and a maximum of 99 images in a class.
  • CDDB-Hard Li et al. (2023) [^3] https://coral79.github.io/CDDB_web/. As a dataset mixed with real-world and model-generated images, the continual deepfake detection benchmark (CDDB) aims to simulate real-world deepfakes’ evolution. The authors put out three scenarios for evaluation, ‘EASY,’ ‘Hard,’ and ’Long,’ respectively. Following Wang et al. (2022b), we choose the ’Hard’ track for evaluation since it poses a more challenging problem due to its complexity.
  • CORe50 Lomonaco & Maltoni (2017) [^4] https://vlomonaco.github.io/core50/index.html#dataset. Composed of 11 sessions characterized by different backgrounds and lighting, CORe50 is built for continual object recognition. Numerous RGB-D images are divided into 8 indoor sessions for training and 3 outdoor sessions for testing, and each session includes a sequence of about 300 frames for all 50 objects.
B.2 Domain Sequences
@@
Appendix C Compared Methods
@@
  • Replay Ratcliff (1990) is an exemplar-based method, which saves a set of exemplars from previous domains (i.e., in this paper, we save 10 exemplars per class) and replay them when learning new domains. Hence, forgetting can be alleviated since the model can revisit informative instances from previous domains when learning new ones. Of note, for classes with fewer than 10 instances, repeatable sampling is allowed to conform to the requirement.
@@
  • iCaRL Rebuffi et al. (2017) is a knowledge distillation-based continual learning algorithm, which saves the previous model in memory. During updating, apart from the cross-entropy loss for learning new tasks, it also introduces the knowledge distillation loss between old and new models to avoid forgetting. It also requires saving an increasing number of exemplars.
@@
  • MEMO Zhou et al. (2023) is an expansion-based continual learning algorithm that partially expands the network to catch new features. As for the implementation, we follow the original paper to decouple the network and expand the last transformer block for each new task.
@@
  • SimpleCIL Zhou et al. (2024a) proposes this simple baseline in pre-trained model-based continual learning. It freezes the backbone representation, extracts the class center of each class, and utilizes a cosine classifier updated by assigning class centers to the classifier weights.
@@
  • L2P Wang et al. (2022d) is the first work introducing prompt tuning in continual learning. With the pre-trained weights frozen, it learns a prompt pool containing many prompts. During training and inference, instance-specific prompts are selected to produce the instance-specific embeddings. However, as alluded to before, learning new domains will lead to the overwriting of existing prompts, thus triggering forgetting.
@@
  • DualPrompt Wang et al. (2022c) extends L2P in two aspects. Apart from the prompt pool and prompt selection mechanism, it further introduces prompts instilled at different depths and task-specific prompts. During training and inference, the instance-specific and task-specific prompts work together to adjust the embeddings.
@@
  • CODA-Prompt Smith et al. (2023) aims to avoid the prompt selection cost in L2P. It treats prompts in the prompt pool as bases and utilizes the attention results to combine multiple prompts as the instance-specific prompt.
@@
  • EASE Zhou et al. (2024b) designs lightweight feature expansion technique with adapters to learn new features as data devolves. To fetch a classifier with the same dimension as ever-expanding features, it utilizes class-wise similarity to complete missing class prototypes.
@@
  • RanPAC McDonnell et al. (2023) extends SimpleCIL by randomly projecting the features into the high-dimensional space and learning the online LDA classifier for final classification.
@@
  • S-iPrompt Wang et al. (2022b) is specially designed for pre-trained model-based domain-incremental learning. It learns task-specific prompts for each domain and saves domain centers in the memory with K-Means. During inference, it first forwards the features to select the nearest domain center via KNN search. Afterward, the selected prompt will be appended to the input.

### v2

@@ III.1 Dataset introduction
  • DomainNet [46]***https://ai.bu.edu/M3SDA/ is a dataset of common objects in six different domains. All domains include 345 categories of objects, such as bracelets, planes, birds, and cellos. The domains include clipart — a collection of clipart images; real — photos and real-world images; sketch — sketches of specific objects; infograph — infographic images with specific objects; painting — artistic depictions of objects in the form of paintings, and quickdraw — drawings of the worldwide players of the game ‘Quick Draw!’. We use the officially recommended version ‘Cleaned’ in this paper.
  • Office-Home [63]https://hemanthdv.github.io/officehome-dataset/ is a benchmark dataset for domain adaptation which contains four domains where each domain consists of 65 categories. The four domains are Art — artistic images in the form of sketches, paintings, ornamentation, etc.; Clipart — a collection of clipart images; Product — images of objects without a background; and Real_World — images of objects captured with a regular camera. It contains 15,500 images, with an average of around 70 images per class and a maximum of 99 images in a class.
  • CDDB-Hard [30]https://coral79.github.io/CDDB_web/. As a dataset mixed with real-world and model-generated images, the continual deepfake detection benchmark (CDDB) aims to simulate real-world deepfakes’ evolution. The authors put out three tracks for evaluation, ‘EASY,’ ‘Hard,’ and ’Long,’ respectively. Following [68], we choose the ’Hard’ track for evaluation since it poses a more challenging problem due to its complexity.
  • CORe50 [36]§§§https://vlomonaco.github.io/core50/index.html#dataset. Composed of 11 sessions characterized by different backgrounds and lighting, CORe50 is built for continual object recognition. Numerous RGB-D images are divided into 8 indoor sessions for training and 3 outdoor sessions for testing, and each session includes a sequence of about 300 frames for all 50 objects.
III.2 Domain Sequences
@@
Table 13: Incremental performances of methods on five representative pre-trained models, i.e., ResNet101, ViT-S/16, ViT-B/16-IN1K, ViT-B/16-IN21K, and ViT-B/16-DINO, comparing the selected best-performing algorithms and Duct. ‘NA’ indicates that these methods are incompatible with ResNet. Duct achieves the optimal performances across all backbones.
ResNet101 ViT-S/16 ViT-B/16-DINO ViT-B/16-IN1K ViT-B/16-IN21K
CODA-Prompt [58] NA 52.90 55.47 58.49 58.99
EASE [86] NA 40.84 38.72 53.04 53.11
S-iPrompt [68] NA 55.12 60.12 58.21 60.13
Duct 50.22 64.55 62.89 67.28 67.37
Table 14: The upper bound on the performance of ViT-B/16-IN21K for all benchmarks, i.e., the performance achieved by fine-tuning on the union of the entire dataset.
CDDB-Hard Core50 DomainNet Office-Home
Upperbound 93.44 96.29 76.28 87.92
Appendix IV Compared Methods
@@
  • Replay [51] is an exemplar-based method, which saves a set of exemplars from previous domains (i.e., in this paper, we save 10 exemplars per class) and replay them when learning new domains. Hence, forgetting can be alleviated since the model can revisit informative instances from previous domains when learning new ones. Of note, for classes with fewer than 10 instances, repeatable sampling is allowed to conform to the requirement.
@@
  • iCaRL [52] is a knowledge distillation-based continual learning algorithm, which saves the previous model in memory. During updating, apart from the cross-entropy loss for learning new tasks, it also introduces the knowledge distillation loss between old and new models to avoid forgetting. It also requires saving an increasing number of exemplars.
@@
  • MEMO [85] is an expansion-based continual learning algorithm that partially expands the network to catch new features. As for the implementation, we follow the original paper to decouple the network and expand the last transformer block for each new task.
@@
  • SimpleCIL [88] proposes this simple baseline in pre-trained model-based continual learning. It freezes the backbone representation, extracts the class center of each class, and utilizes a cosine classifier updated by assigning class centers to the classifier weights.
@@
  • L2P [70] is the first work introducing prompt tuning in continual learning. With the pre-trained weights frozen, it learns a prompt pool containing many prompts. During training and inference, instance-specific prompts are selected to produce the instance-specific embeddings. However, as alluded to before, learning new domains will lead to the overwriting of existing prompts, thus triggering forgetting.
@@
  • DualPrompt [69] extends L2P in two aspects. Apart from the prompt pool and prompt selection mechanism, it further introduces prompts instilled at different depths and task-specific prompts. During training and inference, the instance-specific and task-specific prompts work together to adjust the embeddings.
@@
  • CODA-Prompt [58] aims to avoid the prompt selection cost in L2P. It treats prompts in the prompt pool as bases and utilizes the attention results to combine multiple prompts as the instance-specific prompt.
@@
  • EASE [86] designs lightweight feature expansion technique with adapters to learn new features as data devolves. To fetch a classifier with the same dimension as ever-expanding features, it utilizes class-wise similarity to complete missing class prototypes.
@@
  • RanPAC [40] extends SimpleCIL by randomly projecting the features into the high-dimensional space and learning the online LDA classifier for final classification.
@@
  • S-iPrompt [68] is specially designed for pre-trained model-based domain-incremental learning. It learns task-specific prompts for each domain and saves domain centers in the memory with K-Means. During inference, it first forwards the features to select the nearest domain center via KNN search. Afterward, the selected prompt will be appended to the input.
Appendix V More Experiments
To further demonstrate the effectiveness of Duct in different scenarios, we conduct a series of more comprehensive experiments in this section.
V.1 More backbones
Apart from ViT-B/16-IN1K and ViT-B/16-IN21K, we also evaluate additional backbones at different scales, including ResNet101 [22], ViT-S/16, and ViT-B/16-DINO [11]. The results are reported in Table 13. Note that the compared methods are all based on prompt tuning technique [28] which is designed for Vision Transformers [17]. Hence, these compared methods are incompatible with ResNet structures, while Duct can still boost ResNet in continual learning tasks, further highlighting that Duct is a general and versatile framework. The outcomes demonstrate that Duct outperforms all competitors across all PTMs.
V.2 Upper bound performance
To investigate the upper-bound performance of each dataset, we conduct joint training and present the results in Table 14. We adopt a learning rate of 0.0001 for the representation layer and 0.01 for the classification layer, with a batch size of 128. We report the optimal results in the table.
V.3 Extending Duct to Class-Incremental Learning
Since Duct is a general framework, it can also be applied to the class-incremental learning (CIL) scenario. By equally splitting 345 classes of DomainNet into 5 incremental stages, we formulate a class-incremental learning setting, each containing 69 classes. We keep the other settings the same as in the main paper and report the results in Table 15. The results indicate that DUCT can also be applied to the class-incremental learning setting, which we shall explore in future work.
Table 15: Incremental performances of selected methods on CIL setting, evaluated on the most challenging benchmark dataset, DomainNet, using ViT-B/16-IN21K as the backbone. ‘B0’ indicates that classes are uniformly split across tasks. The results indicate that Duct performs effectively in CIL scenarios.
Method DomainNet B0 Inc69
$\bar{\mathcal{A}}$ ${\mathcal{A}_{B}}$
CODA-Prompt [58] 72.50 71.05
EASE [86] 69.29 68.45
S-iPrompt [68] 73.79 71.81
Duct 75.29 74.68